Sam is 10 years older than Jack. Twice Sam's age subtracted from 6 times Jack's age is 7 more than Jack's age. Find Jack and Sam's age.
Answer by ankor@dixie-net.com(22740)

Sam is 10 years older than Jack.
S = J+10
:
Twice Sam's age subtracted from 6 times Jack's age is 7 more than Jack's age.
6J - 2S = J + 7
6J - J - 2S = 7
5J - 2S = 7
:
Substitute (J+10) for S in the above equation:
5J - 2(J+10) = 7
5J - 2J - 20 = 7
5J - 2J = 7 + 20
3J = 27
J = 27/3
J = 9 yrs old, which makes S = 19 yrs old
:
:
Check our solution in the statement:
"Twice Sam's age subtracted from 6 times Jack's age is 7 more than Jack's age."
6(9) - 2(19) = 9 + 7
54 - 38 = 16
:
:
Does this make sense to you?